Elimination and substitution in solving a system of equation.
3x -2y = 7
2x + y =7

Answer by mananth(16946)

3x -2y = 7.............1
2x + y =7..............2
..
multiply equation 2 by 2 and it to equation 1
2(2x+y)+(3x-2y)=14+7
4x+2y+3x-2y=21
7x=21
divide both sides by 7
x=3
plug value of x in equation 3x-3y=7
3*3-2y=7
9-2y=7
add-9 to both sides
9-9-2y=-9+7
-2y=-2
y=1